# Provable Benefit of Annealed Langevin Monte Carlo for Non-log-concave Sampling

### v1

@@ Appendix C Proof of Example 1
By definition of W2 distance, we have
$W_{2}^{2}(\widehat{\pi}_{\lambda},\widehat{\pi}_{\lambda+\delta})$ $\leq\operatorname{\mathbb{E}}\left[\left\|X-Y\right\|^{2}\right]$
$=\operatorname{\mathbb{E}}_{I}\left[\operatorname{\mathbb{E}}_{\eta}\left[\left\|\left(\frac{\beta}{\lambda+\beta}-\frac{\beta}{\lambda+\delta+\beta}\right)y_{I}+\left(\frac{1}{\sqrt{\lambda+\beta}}-\frac{1}{\sqrt{\lambda+\delta+\beta}}\right)\eta\right\|^{2}\right]\right]$
$=\operatorname{\mathbb{E}}_{I}\left[{\left(\frac{\beta}{\lambda+\beta}-\frac{\beta}{\lambda+\delta+\beta}\right)^{2}\left\|y_{I}\right\|^{2}+\left(\frac{1}{\sqrt{\lambda+\beta}}-\frac{1}{\sqrt{\lambda+\delta+\beta}}\right)^{2}d}\right]$
@@
$\left|\dot{\widehat{\pi}}\right|^{2}_{\lambda}=\lim_{\delta\to 0}\frac{W_{2}^{2}(\widehat{\pi}_{\lambda},\widehat{\pi}_{\lambda+\delta})}{\delta^{2}}\leq\frac{\beta^{2}r^{2}}{(\lambda+\beta)^{4}}+\frac{d}{4(\lambda+\beta)^{3}}.$
By time reparameterization $\pi_{\theta}=\widehat{\pi}_{\lambda(\theta)}$, $\left|\dot{\pi}\right|_{\theta}=\left|\dot{\widehat{\pi}}\right|_{\lambda(\theta)}\left|\dot{\lambda}(\theta)\right|$. With $\lambda(\theta)=\lambda_{0}(1-\theta)^{\gamma}$, $\left|\dot{\lambda}(\theta)\right|=\left|\gamma\lambda_{0}(1-\theta)^{\gamma-1}\right|\leq\gamma\lambda_{0}\lesssim\lambda_{0}$. Therefore,
${\cal A}_{\eta,\lambda}(\pi)$ $=\int_{0}^{1}\left|\dot{\pi}\right|_{\theta}^{2}{\rm d}\theta=\int_{0}^{1}\left|\dot{\widehat{\pi}}\right|_{\lambda(\theta)}^{2}\left|\dot{\lambda}(\theta)\right|^{2}{\rm d}\theta$
$\lesssim\lambda_{0}\int_{0}^{1}\left|\dot{\widehat{\pi}}\right|_{\lambda(\theta)}^{2}\left|\dot{\lambda}(\theta)\right|{\rm d}\theta=\lambda_{0}\int_{0}^{\lambda_{0}}\left|\dot{\widehat{\pi}}\right|_{\lambda}^{2}{\rm d}\lambda$
$=\lambda_{0}\int_{0}^{\lambda_{0}}\left(\frac{\beta^{2}r^{2}}{(\lambda+\beta)^{4}}+\frac{d}{4(\lambda+\beta)^{3}}\right){\rm d}\lambda$
@@
It follows from the proof that as long as $\max_{\theta\in[0,1]}\left|\dot{\lambda}(\theta)\right|$ is bounded by a polynomial function of $\beta$, $d$ and $r$, so is the action ${\cal A}_{\eta,\lambda}(\pi)$.
Appendix D Supplementary Lemmas
Lemma 4 ([chewi2022log, Lemma 4.E.1]).
Consider a probability measure $\mu\propto{\rm e}^{-U}$ on $\mathbb{R}^{d}$. If $\nabla^{2}U\preceq\beta I$ for some $\beta>0$, then $\operatorname{\mathbb{E}}_{\mu}\left[\left\|\nabla U\right\|^{2}\right]\leq\beta d$.
@@ Appendix D Supplementary Lemmas
Under Assumption 1, $\pi_{\theta}$ defined in Equation 3 has finite second-order moment when $\eta(\theta)\in[0,1]$.
@@
When $\eta(\theta)=1$, $\pi_{\theta}\propto\exp\left(-V-\frac{\lambda(\theta)}{2}\left\|\cdot\right\|^{2}\right)\leq\exp\left(-V\right)$, and the claim is straightforward. Otherwise, by convexity of $u\mapsto{\rm e}^{-u}$, we have
$\exp\left(-\eta(\theta)V-\frac{\lambda(\theta)}{2}\left\|\cdot\right\|^{2}\right)$ $=\exp\left(-\eta(\theta)V-(1-\eta(\theta))\frac{\lambda(\theta)}{2(1-\eta(\theta))}\left\|\cdot\right\|^{2}\right)$
$\leq\eta(\theta)\exp\left(-V\right)+(1-\eta(\theta))\exp\left(-\frac{\lambda(\theta)}{2(1-\eta(\theta))}\left\|\cdot\right\|^{2}\right).$
Multiplying both sides by $\left\|\cdot\right\|^{2}$ and taking integral over $\mathbb{R}^{d}$, we see that $\operatorname{\mathbb{E}}_{\pi_{\theta}}\left[\left\|\cdot\right\|^{2}\right]<{+\infty}$. ∎

### v2

@@ Appendix C Proofs for Annealed LMC
By the definition of the $\text{W}_{\text{2}}$ distance, we have
$W_{2}^{2}(\widehat{\pi}_{\lambda},\widehat{\pi}_{\lambda+\delta})$ $\leq\mathbb{E}\|X-Y\|^{2}$
$=\mathbb{E}_{I}\mathbb{E}_{\eta}\left\|\left(\frac{\beta}{\lambda+\beta}-\frac{\beta}{\lambda+\delta+\beta}\right)y_{I}+\left(\frac{1}{\sqrt{\lambda+\beta}}-\frac{1}{\sqrt{\lambda+\delta+\beta}}\right)\eta\right\|^{2}$
$=\mathbb{E}_{I}\left(\frac{\beta}{\lambda+\beta}-\frac{\beta}{\lambda+\delta+\beta}\right)^{2}\|y_{I}\|^{2}+\left(\frac{1}{\sqrt{\lambda+\beta}}-\frac{1}{\sqrt{\lambda+\delta+\beta}}\right)^{2}d$
@@
$|\dot{\widehat{\pi}}|^{2}_{\lambda}=\lim_{\delta\to 0}\frac{W_{2}^{2}(\widehat{\pi}_{\lambda},\widehat{\pi}_{\lambda+\delta})}{\delta^{2}}\leq\frac{\beta^{2}r^{2}}{(\lambda+\beta)^{4}}+\frac{d}{4(\lambda+\beta)^{3}}.$
By time reparameterization $\pi_{\theta}=\widehat{\pi}_{\lambda(\theta)}$, $|\dot{\pi}|_{\theta}=|\dot{\widehat{\pi}}|_{\lambda(\theta)}|\dot{\lambda}(\theta)|$. With $\lambda(\theta)=\lambda_{0}(1-\theta)^{\gamma}$, $|\dot{\lambda}(\theta)|=\gamma\lambda_{0}(1-\theta)^{\gamma-1}\leq\gamma\lambda_{0}\lesssim\lambda_{0}$. Therefore,
${\cal A}$ $=\int_{0}^{1}|\dot{\pi}|_{\theta}^{2}\mathrm{d}\theta=\int_{0}^{1}|\dot{\widehat{\pi}}|_{\lambda(\theta)}^{2}|\dot{\lambda}(\theta)|^{2}\mathrm{d}\theta$
$\lesssim\lambda_{0}\int_{0}^{1}|\dot{\widehat{\pi}}|_{\lambda(\theta)}^{2}|\dot{\lambda}(\theta)|\mathrm{d}\theta=\lambda_{0}\int_{0}^{\lambda_{0}}|\dot{\widehat{\pi}}|_{\lambda}^{2}\mathrm{d}\lambda$
$=\lambda_{0}\int_{0}^{\lambda_{0}}\left(\frac{\beta^{2}r^{2}}{(\lambda+\beta)^{4}}+\frac{d}{4(\lambda+\beta)^{3}}\right)\mathrm{d}\lambda$
@@
It follows from the proof that as long as $\max_{\theta\in[0,1]}|\dot{\lambda}(\theta)|$ is bounded by a polynomial function of $\beta$, $d$ and $r$, so is the action ${\cal A}$. $\square$
Appendix E Supplementary Lemmas
Lemma 6 (Chewi (2024, Lemma 4.E.1)).
Consider a probability measure $\mu\propto{\rm e}^{-U}$ on $\mathbb{R}^{d}$. If $\nabla^{2}U\preceq\beta I$ for some $\beta>0$, then $\mathbb{E}_{\mu}{\|\nabla U\|^{2}}\leq\beta d$.
Lemma 7 (Cheng et al. (2023, Lemma 4)).
For a Gaussian mixture distribution $\pi=\sum_{i}w_{i}\operatorname{{\cal N}}\left(\mu_{i},\sigma^{2}I\right)$, we have
$\left(\frac{1}{\sigma^{2}}-\frac{\max_{i,j}\|\mu_{i}-\mu_{j}\|^{2}}{2\sigma^{4}}\right)I\preceq-\nabla^{2}\log\pi\preceq\frac{1}{\sigma^{2}}I.$
Remark.
This is a slightly improved version, as in the original lemma the authors omitted the factor $2$ in the denominator on the l.h.s.
@@ Appendix E Supplementary Lemmas
Under Assump. 2, $\pi_{\theta}$ defined in Eq. 5 has finite second-order moment when $\eta(\theta)\in[0,1]$.
@@
When $\eta(\theta)=1$, $\pi_{\theta}\propto\exp\left(-V-\frac{\lambda(\theta)}{2}\|\cdot\|^{2}\right)\leq\exp\left(-V\right)$, and the claim is straightforward. Otherwise, by convexity of $u\mapsto{\rm e}^{-u}$, we have
$\exp\left(-\eta(\theta)V-\frac{\lambda(\theta)}{2}\|\cdot\|^{2}\right)$ $=\exp\left(-\eta(\theta)V-(1-\eta(\theta))\frac{\lambda(\theta)}{2(1-\eta(\theta))}\|\cdot\|^{2}\right)$
$\leq\eta(\theta)\exp\left(-V\right)+(1-\eta(\theta))\exp\left(-\frac{\lambda(\theta)}{2(1-\eta(\theta))}\|\cdot\|^{2}\right).$
Multiplying both sides by $\|\cdot\|^{2}$ and taking integral over $\mathbb{R}^{d}$, we see that $\mathbb{E}_{\pi_{\theta}}\|\cdot\|^{2}<{+\infty}$. ∎
Appendix F Further Details of Experiments in Sec. 6
Hyperparameter settings.
For all the experiments, we use the annealing schedule $\lambda(\theta)=5(1-\theta)^{10}$ and $\eta(\theta)\equiv 1$. The step size for ALMC is designed to follow a quadratic schedule: the step size at the $\ell$-th iteration (out of $M$ total iterations, $\ell\in\left[\!\left[1,M\right]\!\right]$) is given by $-\frac{s_{\max}-s_{\min}}{M^{2}/4}\left(\ell-\frac{M}{2}\right)^{2}+s_{\max}$, which increases on $\left[0,\frac{M}{2}\right]$ and decreases on $\left[\frac{M}{2},M\right]$, with a maximum of $s_{\max}$ and a minimum of $s_{\min}$. For $r=2,5,10,15,20,25,30$, the $M$ we choose are $200,500,2500,10000,20000,40000,60000$, respectively, and we set $s_{\max}=0.05$ and $s_{\min}=0.01$, which achieve the best performance across all settings after tuning.
KL divergence estimation.
We use the Information Theoretical Estimators (ITE) toolbox (Szabó, 2014) to empirically estimate the KL divergence. In all cases, we generate $1000$ samples using ALMC, and an additional $1000$ samples from the target distribution. The KL divergence is estimated using the ite.cost.BDKL_KnnK() function, which leverages $k$-nearest-neighbor techniques to compute the divergence from the target distribution to the sampled distribution.
Regression coefficients.
We also compute the linear regression approximation of the curves in Fig. 2 via sklearn.linear_model.LinearRegression. The blue curve has a slope of $2.841$ and an intercept of $1.257$, while the orange one has a slope of $2.890$ and an intercept of $0.904$. The $R^{2}$ scores, calculated using sklearn.metrics.r2_score, are $0.995$ and $0.997$, respectively.